A business owner spent $32,000 to upgrade her business showroom she paid 15% as a down payment and financed the balance of the purchase with a 24-month fixed installment loan with an APR of 3% determined the business owners total finance charge and monthly payment.
A total finance charge=1066.24; monthly payment =1177.76
b. total finance charge=4033.76;monthly payment =1301.41
c.total finance charge=856.80;monthly payment =1169.03
d.total finance charge=3917.44;monthly payment =1498.93

Answer by Theo(13342)

the total cost was 32,000
15% was paid up front = .15 * 32,000 = 4,800.
amount to be financed is 32,000 - 4,800 = 27,200.

present value of loan is 27,200
interest rate of loan is 3% per year / 12 = .25% per month.
number of months is 24.
payment is made at the end of each month.
future value is set to 0.

make your inputs and select PMT and the calculator gives you what the monthly payment is.

i used the calculator at https://arachnoid.com/finance/

the result looks like this:

$$$

you enter the present value as 27,200 because that's money you got from the bank up front.

enter as 27200 (no commas).

the calculator will give you the payment as a negative value because that's money you are paying the bank.

selection c looks like the answer.

selection c is: total finance charge = 856.80; monthly payment = 1169.03.

monthly payment is 1169.03.

24 months of payments equals a total layout of 24 * 1169.09 = 28,058.16.

finance charge is 28,058.16 minus 27,200 = 858.16.

the monthly payment matches selection c, but the finance charge is a little off, with a difference of 858.16 minus 856.80 = 1.36.

that can't be explained by rounding, but selection c should be your selection anyway since the payment is correct and nothing else comes close.

regardless, selection c is the correct answer.

none of the other selections are even close.